# Nicholas J. Pritchard International Centre for Radio Astronomy Research

### v1

@@ III-B4 Step-Forward
(a) Latency encoded example.
(b) Rate encoded example.
(c) Delta encoded example.
(d) Step-Forward encoded example with ‘first’ exposure mode.
(e) Step-Forward encoded example with ‘direct’ exposure mode.
(f) Step-Forward encoded example with ‘latency’ exposure mode.
(g) Original input spectrogram and associated RFI mask.
Figure 1: Example spike raster plots for all explored encodings, original spectrogram, and associated RFI mask.
@@ III-C Training Strategy
We train a simple feedforward SNN on a dataset from the Hydrogen Epoch of Reionization Array (HERA) simulator [11]. This dataset, described fully by Mesarcik et al. [9], comprises 420 training spectrograms, 140 test spectrograms, and 2.76% RFI contamination by pixel count. Each spectrogram covers a 30-minute simulation integrating every 3.52 seconds (producing 512 time-steps) and 512 frequency channels from 105MHz to 195MHz. The RFI is synthetic but modeled from satellite communications (localized in frequency, covering all time), lightning (covering all frequencies, localized in time), ground communication (localized in frequency, partially localized in time), and impulse blips (localized in frequency and time). After encoding into spikes, we split each spectrogram further into $32\times 32$ sub-spectrogram patches, allowing for a smaller test model and less intensive training requirements. We chose to train a small two-layer fully connected network outlined in Table I to focus on the effectiveness of each encoding method specifically.
@@ III-D Hyper-Parameter Tuning
We employ Optuna [16] for hyper-parameter optimization, specifically the tree-structured Parzan estimation algorithm, to simultaneously conduct multi-variate optimization for AUROC, AUPRC, MSE, and F1-Score. Table II contains the bounds of each variable included in the optimization for all encoding methods. We conduct 100 trials for each encoding method.
@@
For each encoding method, we take the trials with the best score in any metric into a final set and subsequently compare these trials between all encoding techniques to determine the overall best encoding method and set of parameters. We discuss the result of this process in Section IV as they comprise an integral part of our investigation. Finally, we perform fifty repeat trials for the best parameters in each encoding method to confirm their performance.
@@ IV-A Encoding Comparison
We present the results of our hyper-parameter searches in Table III. No prior investigation has been conducted into encoding radio astronomy visibility data into SNNs for RFI detection, and as such, these results are of interest. For each method, we present the number of trials conducted, the best candidate set (an artifact of performing a multi-variate optimization), and the best trials whose performance in at least one metric (accuracy, MSE, AUROC, AUPRC, or F1-Score) is best. We ran fifty trials for the best parameters for each encoding method and present the results in Table IV.
TABLE III: Hyper-Parameter search using Optuna multi-variate optimization. The best trial for each method is bolded, and the best SNN encoding method is underlined. A standard ANN narrowly produced the highest overall performing hyper-parameters in a single trial, producing top AUPRC and F1-Scores.
Encoding Method Accuracy MSE AUROC AUPRC F1-Score Batch Size Epochs Beta Exposure
@@
Encoding Method Accuracy MSE AUROC AUPRC F1 Num Trials
Delta 0.963 0.037 0.037 0.037 0.840 0.013 0.630 0.075 0.571 0.126 50
Latency 0.988 0.004 0.012 0.004 0.929 0.004 0.785 0.035 0.761 0.050 50
Step-Forward-Direct 0.976 0.015 0.024 0.015 0.856 0.022 0.681 0.071 0.643 0.105 50
Step-Forward-First 0.975 0.015 0.025 0.015 0.859 0.021 0.679 0.074 0.637 0.109 50
Step-Forward-Latency 0.972 0.015 0.028 0.015 0.863 0.021 0.662 0.073 0.611 0.108 50
Rate 0.943 0.009 0.057 0.009 0.513 0.003 0.066 0.006 0.059 0.002 50
ANN 0.816 0.181 0.184 0.181 0.846 0.081 0.595 0.110 0.368 0.223 50
The best SNN encoding scheme based on the hyper-parameter search is the step-forward-first method, with an accuracy of 99%, near-zero MSE of $0.005$, AUROC of $0.89$, AUPRC of $0.857$, and F1-Score of $0.844$. The step-forward-direct and step-forward-latency methods also perform strongly. Combining the delta-modulated changes and sustained value helps encode visibility data, which is similarly effective to the speech2spikes encoding method, albeit demonstrated on a classification task. Latency and Delta-modulation encoding perform similarly well to each other. Rate encoding is deeply non-competitive, with an accuracy of 15% and an F1-score of $0.064$. We conducted a hyper-parameter search and evaluation with an ANN of the identical two-layer architecture of the SNNs. We found that the single-trial performance in the hyper-parameter search narrowly outperformed all SNN methods by providing top AUPRC and F1-Scores.

### v2

@@ III-B4 Step-Forward
(a) Latency encoded example. The background gradient is reflected across the entire spike-train, and the banding and blips are easily visible.
(b) Rate encoded example. The features are not comprehensible in this encoding for such a small exposure value.
(c) Delta encoded example. While most features are legible, and removing the gradient may be helpful, this encoding misses the banding below the background as a sub-threshold feature.
(d) Step-Forward encoded example with ‘first’ exposure mode. This encoding clearly shows the banding effect, however, the blips are not easily identifiable.
(e) Step-Forward encoded example with ‘direct’ exposure mode. This encoding is highly active however the blips are still hidden.
(f) Step-Forward encoded example with ‘latency’ exposure mode. This encoding, much like the original latency encoding, makes the background gradient, banded RFI, and blips easily identifiable.
(g) The original input spectrogram and associated RFI mask. This example gives a good indication of the type of challenge RFI detection presents. This example contains banded RFI present across all frequencies that is both above and below the background environment, two ‘blips’ tightly localized in frequency and time, and a gradient in the background.
Figure 1: Example spike raster plots for all explored encodings, the original spectrogram, and associated RFI mask. Encodings are exposed for four time-steps where applicable.
@@ III-C Training Strategy
We train a simple feedforward SNN on a dataset from the Hydrogen Epoch of Reionization Array (HERA) simulator [11]. This dataset, described fully by Mesarcik et al. [9], comprises 420 training spectrograms, 140 test spectrograms, and 2.76% RFI contamination by pixel count. Each spectrogram covers a 30-minute simulation integrating every 3.52 seconds (producing 512 time-steps) and 512 frequency channels from 105MHz to 195MHz. The RFI is synthetic but modeled from satellite communications (localized in frequency, covering all time), lightning (covering all frequencies, localized in time), ground communication (localized in frequency, partially localized in time), and impulse blips (localized in frequency and time). After encoding into spikes, we split each spectrogram further into $32\times 32$ sub-spectrogram patches, allowing for a smaller test model and less intensive training requirements. We chose to train a small two-layer fully connected network outlined in Table I, avoiding complex structures and recurrent connects to simplify training and focus on the effectiveness of each encoding method specifically, even if this leads to lower overall task performance.
@@ III-D Hyper-Parameter Tuning
We employ Optuna [16] for hyper-parameter optimization, specifically the tree-structured Parzan estimation algorithm, to simultaneously conduct multi-variate optimization for Area Under the Receiver Operating Characteristic curve (AUROC), Area Under the Precision-Recall Curve (AUPRC), Accuracy, and F1-Score. AUROC evaluates the ratio of True Positive Rate (TPR) and False Positive Rate (FPR) across all possible classification thresholds. AUPRC gives the ratio of precision and recall across several thresholds, in this case referring to the proportion of correctly classified RFI over all RFI predictions. Recall in this case is the TPR. Accuracy is the per-pixel output accuracy. For this synthetic radio astronomy dataset, we expect accuracy to be very high as true RFI flags are very sparse; a completely silent network would still be more than 90% accurate. F1-Score is the harmonic mean of precision and recall at a given threshold. The sparsity of RFI in this dataset means we generally expect high AUROC and high accuracy values with lower AUPRC and F1 scores. A good-performing model needs to be silent the majority of the time while providing confident RFI detections, hence our inclusion of both raw accuracy and f1-scores. Table II contains the bounds of each variable included in the optimization for all encoding methods. We conduct 100 trials for each encoding method.
@@
For each encoding method, we take the trials with the best score in any metric into a final set and subsequently compare these trials between all encoding techniques to determine the overall best encoding method and set of parameters. Balancing the rate of neuron-state decay (beta) with the exposure of each timestep in the spectrogram is challenging; more exposure increases the fidelity of the input data but dilutes the information propagating over real time. Moreover, a changing exposure changes the input data shape hence why we allow Optuna to select batch size. We discuss the result of this process in Section IV as they comprise an integral part of our investigation. Finally, we perform fifty repeat trials for the best parameters in each encoding method to confirm their performance.
@@ IV-A Encoding Comparison
We present the results of our hyper-parameter searches in Table III. No prior investigation has been conducted into encoding radio astronomy visibility data into SNNs for RFI detection, and as such, these results are of interest. For each method, we present the number of trials conducted, the best candidate set (an artifact of performing a multi-variate optimization), and the best trials whose performance in at least one metric (accuracy, AUROC, AUPRC, or F1-Score) is best. We ran fifty trials for the best parameters for each encoding method and present the results in Table IV.
TABLE III: Hyper-Parameter search using Optuna multi-variate optimization. We show the top three to five trials per encoding method, as Optuna selects during its multi-objective optimization routine. A range of beta and exposure values produce decent results in all cases. Trials with smaller exposure tend to score higher in raw accuracy. Larger exposures tend to combine with larger beta values to allow information to propagate throughout the entire spectrogram sample. The sparsity of RFI in radio astronomy spectrograms means models tend to perform better in accuracy and AUROC than in AUPRC and F1-Scores, where performance focuses on the positive class. The best trial for each method is bolded, and the best SNN encoding method is underlined. A standard ANN narrowly produced the highest overall performing hyper-parameters in a single trial, producing top AUPRC and F1-Scores.
Encoding Method Accuracy AUROC AUPRC F1-Score Batch Size Epochs Beta Exposure
@@
Encoding Method Accuracy AUROC AUPRC F1 Num Trials
Delta 0.963 0.037 0.840 0.013 0.630 0.075 0.571 0.126 50
Latency 0.988 0.004 0.929 0.004 0.785 0.035 0.761 0.050 50
Step-Forward-Direct 0.976 0.015 0.856 0.022 0.681 0.071 0.643 0.105 50
Step-Forward-First 0.975 0.015 0.859 0.021 0.679 0.074 0.637 0.109 50
Step-Forward-Latency 0.972 0.015 0.863 0.021 0.662 0.073 0.611 0.108 50
Rate 0.943 0.009 0.513 0.003 0.066 0.006 0.059 0.002 50
ANN 0.816 0.181 0.846 0.081 0.595 0.110 0.368 0.223 50
The best SNN encoding scheme based on the hyper-parameter search is the step-forward-first method, with an accuracy of 99%, AUROC of $0.89$, AUPRC of $0.857$, and F1-Score of $0.844$. The step-forward-direct and step-forward-latency methods also perform strongly. Combining the delta-modulated changes and sustained value helps encode visibility data, which is similarly effective to the speech2spikes encoding method, albeit demonstrated on a classification task. Latency and Delta-modulation encoding perform similarly well to each other. Rate encoding is deeply non-competitive, with an accuracy of 15% and an F1-score of $0.064$. We conducted a hyper-parameter search and evaluation with an ANN of the identical two-layer architecture of the SNNs. We found that the single-trial performance in the hyper-parameter search narrowly outperformed all SNN methods by providing top AUPRC and F1-Scores.